# d3LLM: Ultra-Fast dLLM using Pseudo-Trajectory Distillation

### v1

@@ 5 Conclusion
In this paper, we observe a fundamental accuracy–parallelism trade-off in diffusion large language models (dLLMs). However, existing methods typically focus on only one side of the coin, targeting either efficiency or performance. Moreover, previous training approaches with random masking provide no guidance on which tokens can be safely decoded early, while previous multi-block decoding methods degrade quality due to incomplete or erroneous context from preceding blocks. To address these limitations, we propose d3LLM (pseuDo-Distilled-Diffusion LLM) framework, striking a balance between the accuracy and the parallelism: at training time, pseudo-trajectory distillation teaches the model which tokens can be confidently decoded early by leveraging the teacher’s unmasking order; at inference time, entropy-based multi-block decoding with a KV-cache refresh mechanism enables high parallelism while maintaining accuracy. To better evaluate dLLMs, we also introduce AUP (Accuracy Under Parallelism), a new metric that jointly measures accuracy and parallelism. Extensive experiments on LLaDA, Dream, and Dream-Coder, demonstrate that our d3LLM framework achieves the highest AUP score on 9 out of 10 benchmark tasks, delivering up to $3.6\times$$5\times$ speedup over AR models (Qwen-2.5-7B-it) depending on the GPU platform, and $10\times$ speedup over vanilla LLaDA/Dream, with negligible accuracy degradation.
@@ A.3 Coder Models
A.4 Details of Contenders
@@ A.5 Details of Datasets
A.6 More Implementation Details
@@
Inference Settings.   During inference, we set the maximum generation length to 256 tokens for most tasks and 512 tokens for Dream-Coder experiments. We use greedy decoding with temperature set to 0.0 or 0.1 depending on the specific task. The block size is fixed at 32 tokens for all experiments. For our d3LLM framework with multi-block generation, the entropy threshold ranges from 0.4 to 0.5, the block-add threshold is set to 0.1, and the decoded token threshold is 0.95. The cache delay iteration is typically set to 1–2, depending on the task requirements. For more details on hyperparameter configurations, please refer to our code repository.
A.7 Compare with Speculative Decoding Method
@@ A.8 Further Improvements of d3LLM
Applying to Stronger Foundation dLLMs.   Most recently, there has been rapid progress in converting AR models to dLLMs (Wu et al., 2025a; Cheng et al., 2025; Li et al., 2025a; Fu et al., 2025; Bie et al., 2025). These new methods achieve notably better performance than the earlier foundation dLLMs (LLaDA and Dream) that we use in this work. We claim that, since our d3LLM is a post-training framework consists of a distillation recipe and a decoding strategy that are largely model-agnostic, therefore, it can serve as a plug-and-play component for these stronger dLLMs. Applying our pseudo-trajectory distillation and multi-block decoding strategy to advanced models such as ReFusion (Li et al., 2025a) or LLaDA 2.0 (Bie et al., 2025) may yield further improvements in both accuracy and parallelism. Alternatively, incorporating reinforcement learning techniques (Zhu et al., 2025; Li et al., 2025b; Wang et al., 2025b) may further enhance the effectiveness of d3LLM.
System-Level Optimizations.   All experiments in this work use the HuggingFace Transformers inference backend, which provides a convenient but not fully optimized environment for dLLM inference. Several system-level optimizations could further improve throughput: (i) GPU kernel fusion for the attention and feed-forward operations tailored to the multi-block decoding pattern; (ii) integration with high-performance inference engines such as vLLM, which would require adapting the paged attention mechanism to the bidirectional attention used in dLLMs; and (iii) better memory management strategies for the KV-cache refresh mechanism. We leave these infrastructure improvements to future work.
@@ Appendix B Related Work
With the growing interest of the research community, an increasing number of works have been proposed to improve dLLMs in terms of efficiency or performance. On one hand, acceleration of dLLMs has been an active research area in recent years (Wu et al., 2025b; Ma et al., 2025a; Gao et al., 2025; Ma et al., 2025b; Yang et al., 2025a). Fast-dLLM (Wu et al., 2025b) presents a training-free acceleration method using a block-wise approximate KV-cache mechanism tailored for dLLMs. dKV (Ma et al., 2025a) proposes a delayed KV-cache mechanism that caches key and value states with a delayed and conditioned strategy to accelerate inference. dInfer (Ma et al., 2025b) develops iteration smoothing, hierarchical and credit decoding, and refresh strategies alongside system-level optimizations to accelerate dLLMs across multiple dimensions. Most recently, dParallel (Chen et al., 2025) introduces a learnable parallel decoding mechanism with certainty-forcing loss for distillation (Hinton et al., 2015; Zhou & Jiang, 2004), achieving significant speedup over vanilla dLLMs.

### v2

@@ 5 Conclusion
In this paper, we observe a fundamental accuracy–parallelism trade-off in dLLMs. Existing methods typically focus on only one side of the coin, targeting either efficiency or performance. Moreover, previous training approaches with random masking provide no guidance on which tokens can be safely decoded early, while previous multi-block decoding methods degrade quality due to incomplete or erroneous context. To this end, we propose d3LLM (pseuDo-Distilled-Diffusion LLM) framework, striking a balance between the accuracy and the parallelism: at training time, pseudo-trajectory distillation teaches the model which tokens can be confidently decoded early by leveraging the teacher’s unmasking order; at inference time, entropy-based multi-block decoding with a KV-cache refresh mechanism enables high parallelism while maintaining accuracy. To better evaluate dLLMs, we also introduce AUP (Accuracy Under Parallelism), a new metric that jointly measures accuracy and parallelism. Experiments on LLaDA, Dream, and Dream-Coder, demonstrate that d3LLM achieves the highest AUP score on 9 out of 10 benchmark tasks, delivering up to $3.6\times$$5\times$ speedup over AR models depending on the GPU platform, and $10\times$ speedup over vanilla LLaDA/Dream, with negligible accuracy degradation.
Acknowledgments
The authors would like to thank Chongxuan Li for the insightful and helpful discussions.
@@ A.3 Coder Models
A.4 Validation of the Hyperparameter in AUP Metric
In this section, we validate different choices of hyperparameter $\alpha$ as defined in the AUP (Accuracy Under Parallelism) metric. Recall that $\alpha$ controls the penalty for accuracy degradation: a larger $\alpha$ increases sensitivity to performance drops, causing the contribution of throughput to decay exponentially with the error rate. We evaluate all methods on GSM8K-CoT with $\alpha\in\{1,2,3,5,10\}$ and report the AUP scores in Tables 10 and 10.
As shown in the tables, d3LLM consistently achieves the highest AUP scores across different $\alpha$ values for both LLaDA-based and Dream-based methods. When $\alpha$ is small (e.g., $\alpha=1$), the metric is more tolerant to accuracy drops and primarily rewards parallelism gains. As $\alpha$ increases, the penalty for accuracy degradation becomes more severe, and methods with larger accuracy drops see their AUP scores decrease more sharply. Notably, our d3LLM maintains competitive AUP scores under all $\alpha$ values, demonstrating that our approach achieves a favorable balance between parallelism improvement and accuracy preservation. We set $\alpha=3$ as the default in our main experiments, as it provides a reasonable trade-off between accuracy and parallelism.
Remark 3 (Why Not Optimize AUP Directly?).
Although AUP effectively captures the accuracy–parallelism trade-off, it cannot serve as a training objective because it is non-differentiable. Specifically, the parallelism $\rho_{i}$ measured by TPF is a runtime statistic from the decoding process rather than an explicit function of model parameters $\theta$. Moreover, AUP is defined over a set of parallelism-accuracy pairs $\{(\rho_{i},y_{i})\}_{i=1}^{m}$, each corresponding to a discrete hard confidence threshold selected for decoding, which precludes gradient computation. We therefore treat AUP as an evaluation metric. Instead, inspired by this accuracy–parallelism trade-off, we propose d3LLM as described in Section 3.
Table 9: Sensitivity analysis of $\alpha$ on LLaDA-based methods. We report AUP scores with different $\alpha$ values on GSM8K-CoT.
Method AUP Score $\uparrow$
$\alpha=1$ $\alpha=2$ $\alpha=3$ $\alpha=5$ $\alpha=10$
Qwen-2.5-7B-it 74.1 74.1 74.1 74.1 74.1
LLaDA 72.6 72.6 72.6 72.6 72.6
Fast-dLLM-LLaDA 206.6 206.2 205.8 204.9 202.8
D2F-LLaDA 214.8 214.3 213.8 212.7 210.1
dParallel-LLaDA 370.9 364.4 358.1 346.0 318.3
d3LLM-LLaDA 659.4 648.4 637.7 616.8 568.4
Table 10: Sensitivity analysis of $\alpha$ on Dream-based methods. We report AUP scores with different $\alpha$ values on GSM8K-CoT.
Method AUP Score $\uparrow$
$\alpha=1$ $\alpha=2$ $\alpha=3$ $\alpha=5$ $\alpha=10$
Qwen-2.5-7B-it 74.1 74.1 74.1 74.1 74.1
Dream 83.9 83.9 83.9 83.9 83.9
Fast-dLLM-Dream 118.4 117.4 116.5 114.8 111.2
Fast-dLLM-v2 180.1 178.1 176.1 172.3 163.4
dParallel-Dream 249.5 247.6 245.7 242.2 233.8
d3LLM-Dream 402.4 396.8 391.3 380.8 356.8
A.5 Details of Contenders
@@ A.6 Details of Datasets
A.7 More Implementation Details
@@
Extracting Pseudo Trajectory of the Teacher dLLM.   As mentioned in Section 3, we extract the pseudo-trajectory of the teacher dLLM by using the teacher dLLM to generate and record its own decoding trajectory, where we constrain the teacher model to unmask exactly one token at each decoding step, and we continue generation beyond the EOS token so that the output length is exactly $n$. One may observe that for a decoding window $w=\{s,\ldots,s+k\}$, directly using the global trajectory $\mathcal{T}_{s+\lceil kt\rceil}$ to construct the noisy sequence does not guarantee that the mask ratio within the window equals exactly $t$. An alternative approach might be to first sort the tokens within $w$ according to their generation order in the teacher’s trajectory, and then mask the last $\lceil kt\rceil$ tokens to ensure a precise local mask ratio of $t$. However, we argue that this window-specific adaptation is inappropriate: our target is to train a dLLM with a block size of 32, and the teacher’s pseudo-trajectory reflects the generation order under this full block size. Reordering or re-indexing tokens within smaller windows (i.e., $k<32$) would disrupt the global ordering learned by the teacher, potentially introducing inconsistencies between the distillation objective and the actual inference-time decoding behavior. Therefore, we adopt the global trajectory without window-specific modifications, which preserves the teacher’s original generation order.
Inference Settings.   During inference, we set the maximum generation length to 256 tokens for most tasks and 512 tokens for Dream-Coder experiments. We use greedy decoding with temperature set to 0.0 or 0.1, depending on the specific task. The block size is fixed at 32 tokens for all experiments. For our d3LLM framework with multi-block generation, the entropy threshold ranges from 0.4 to 0.5, the block-add threshold is set to 0.1, and the decoded token threshold is 0.95. The cache delay iteration is typically set to 1–2, depending on the task requirements. For more details on hyperparameter configurations, please refer to our code repository.
A.8 Compare with Speculative Decoding Method
@@ A.9 Further Improvements of d3LLM
Applying to Stronger Foundation dLLMs.   Most recently, there has been rapid progress in converting AR models to dLLMs (Wu et al., 2025a; Cheng et al., 2025; Li et al., 2025a; Fu et al., 2025; Bie et al., 2025). These new methods achieve notably better performance than the earlier foundation dLLMs (LLaDA and Dream) that we use in this work. We claim that, since our d3LLM is a post-training framework that consists of a distillation recipe and a decoding strategy that are largely model-agnostic, it can serve as a plug-and-play component for these stronger dLLMs. Applying our pseudo-trajectory distillation and multi-block decoding strategy to advanced models such as ReFusion (Li et al., 2025a) or LLaDA 2.0 (Bie et al., 2025) may yield further improvements in both accuracy and parallelism. Alternatively, incorporating reinforcement learning techniques (Zhu et al., 2025; Li et al., 2025b; Wang et al., 2025b) and quantization techniques (Lin et al., 2025) may further enhance the effectiveness of d3LLM.
System-Level Optimization.   All experiments in this work use the HuggingFace Transformers inference backend, which provides a convenient but not fully optimized environment for dLLM inference. Several system-level optimizations could further improve throughput: (i) GPU kernel fusion for the attention and feed-forward operations tailored to the multi-block decoding pattern; (ii) integration with high-performance inference engines such as vLLM, which would require adapting the paged attention mechanism to the bidirectional attention used in dLLMs; and (iii) better memory management strategies for the KV-cache refresh mechanism. We leave these infrastructure improvements to future work.
@@ Appendix B Related Work
With the growing interest of the research community, an increasing number of works have been proposed to improve dLLMs in terms of efficiency or performance. On one hand, acceleration of dLLMs has been an active research area in recent years (Wu et al., 2025b; Ma et al., 2025a; Gao et al., 2025; Ma et al., 2025b; Yang et al., 2025a; Lin et al., 2026). Fast-dLLM (Wu et al., 2025b) presents a training-free acceleration method using a block-wise approximate KV-cache mechanism tailored for dLLMs. dKV (Ma et al., 2025a) proposes a delayed KV-cache mechanism that caches key and value states with a delayed and conditioned strategy to accelerate inference. dInfer (Ma et al., 2025b) develops iteration smoothing, hierarchical and credit decoding, and refresh strategies alongside system-level optimizations to accelerate dLLMs across multiple dimensions. Most recently, dParallel (Chen et al., 2025) introduces a learnable parallel decoding mechanism with certainty-forcing loss for distillation (Hinton et al., 2015; Zhou & Jiang, 2004), achieving significant speedup over vanilla dLLMs.